### v2

@@ Abstract
We study the densest subgraph problem and its variants through the lens of learning-augmented algorithms. We show that, given a reasonably accurate predictor that estimates whether a node belongs to the densest subgraph (e.g., a machine-learning classifier), one can design simple and practical linear-time algorithms that achieve a $(1-\epsilon)$-approximation to the densest subgraph. Our approach also extends to the NP-Hard densest at-most-$k$ subgraph problem and to the directed densest subgraph variant. Finally, we present experimental results demonstrating the effectiveness of our methods.
@@ 1 Introduction
In this work, we study the densest subgraph problem and its variants within the framework of learning-augmented algorithms. This approach is motivated by the observation that real-world data often exhibit underlying patterns that can be captured using machine learning. In this work, we mainly assume access to a predictor (e.g., a machine-learning model) that estimates whether a node belongs to the densest subgraph. Such predictor is typically obtained by training a model on a dataset for which ground-truth labels are available.
@@
Figure 1: Suppose the this is the densest subgraph and the predictor identifies the nodes $y_{1},\ldots,y_{5}$. Returning $y_{1},\ldots,y_{5}$ will yield an infinitely bad solution with density 0 while technically the model has recalled $5/7\approx 71\%$ of the nodes in the densest subgraph.
Exact algorithms for the densest subgraph problem exist using linear programming or maximum flow [15, 28]. For unweighted graphs, Goldberg’s max-flow based algorithm runs in $O(m\log n)$ time. One can also approximate the densest subgraph up to a $1-\epsilon$ factor using the multiplicative weights method to solve the dual linear program followed by a clever rounding algorithm [41, 6, 26] which runs in $O(m\cdot 1/\epsilon^{2}\cdot\log^{2}n)$ time. There is, however, a simple and practical greedy $1/2$-approximation that can be implemented in $O(m)$ time by Charikar [15]; this algorithm keeps removing the node with the lowest degree from the graph and recording the densest subgraph in the process.
One practical issue with the densest subgraph problem is that the densest subgraph is often too large to be useful. Instead, one may consider the problem of finding the densest subgraph with at most $k$ nodes. This variant is NP-Hard. Andersen [2] showed that an $\alpha$-approximation for this problem implies an $O(\alpha^{2})$-approximation for the densest $k$ subgraph problem. Furthermore, under the Exponential Time Hypothesis (ETH), no polynomial-time constant-factor approximation exists for the densest $k$ subgraph problem [32]. Hence, the densest at-most-$k$-subgraph problem is also likely difficult to approximate. For the densest $k$ subgraph problem, the current best polynomial time approximation is $O(n^{1/4})$ [10].
@@
We first address the issue of uniqueness. Throughout, we let $H^{\star}$ denote an arbitrary optimal solution. This assumption is often reasonable in practice: in real-world instances, the optimal solution is typically unique and, moreover, there are usually no competing optimal solutions. That is, any solution whose objective value is close to optimal must substantially overlap with the optimal one. This type of stability phenomenon has been studied extensively in problems such as clustering, independent set, and Max-Cut (see, e.g., [8, 12, 4, 5]).
For the densest subgraph problem, it is possible to enforce uniqueness as follows. If $H_{1}^{\star}$ and $H_{2}^{\star}$ are two solutions to the densest subgraph problem, then $H_{1}^{\star}\cup H_{2}^{\star}$ is also a solution to the densest subgraph problem. To see this, let $\text{OPT}=\textup{d}(H_{1}^{\star})=\textup{d}(H_{2}^{\star})$. Note that $e(H_{1}^{\star})=\text{OPT}|H_{1}^{\star}|$, $e(H_{2}^{\star})=\text{OPT}|H_{2}^{\star}|$, and $e(H_{1}^{\star}\cap H_{2}^{\star})\leq\text{OPT}|H_{1}^{\star}\cap H_{2}^{\star}|$.
@@
Hence, we may enforce uniqueness for the densest subgraph problem by letting $H^{\star}$ be the maximal densest subgraph. The above fact also holds for the densest at-most-$k$-subgraph problem. However, in this variant, $H_{1}^{\star}\cup H_{2}^{\star}$ may not be feasible as its size may exceed $k$.
@@
Intuitively, the two conditions say that the provided partial solution $S$ has most nodes in $H^{\star}$ and few nodes outside $H^{\star}$. In the above definition, one can think of the predictor as a machine learning classifier that predicts whether $v\in H^{\star}$ based on its attributes and $S$ is the set of nodes that are predicted to be in the solution. Note that we can also use two different parameters (e.g., $\epsilon_{1}$ and $\epsilon_{2}$) for the two conditions. We, however, use the same parameter $\epsilon$ for the sake of a clean presentation.
@@
    Given a $(1-\epsilon)$-partial solution $S$ to $H^{\star}$, there exists an $O(m)$-time algorithm that finds a subgraph whose density is at least
@@
    Given an $(\epsilon,\ell_{p})$-probabilistic predictor for $H^{\star}$, for some $p>0$, there exists an $O(m)$-time algorithm that finds a subgraph whose density is at least
@@
The second main result addresses the NP-Hard variant where we want to find the densest subgraph $H^{\star}$ subject to the constraint $|H^{\star}|\leq k$.
@@
    Given a $(1-\epsilon)$-partial solution $S$ to $H^{\star}$, there exists an $O(m)$-time algorithm that finds an at-most-$k$ subgraph whose density is at least
@@
    Given an $(\epsilon,\ell_{p})$-probabilistic predictor for $H^{\star}$, for some $p>0$, there exists an $O(m)$-time algorithm that finds an at-most-$k$ subgraph whose density is at least
@@
We also extend our technique to the directed densest subgraph variant.
@@
Algorithms and their analysis for densest subgraph and densest at-most-$k$ subgraph are presented in Section 2. Section 3 presents simple experiments on the Twitch Ego Nets dataset. In Appendix A, we present the algorithm for densest directed subgraph.
@@ 2.1 Learning-Augmented Algorithms for Densest Subgraph
$H^{*}\setminus S$$H^{*}\cap S$
Figure 2: The approximation guarantee comes from the observation that the number of edges induced by $H^{\star}\setminus S$ is small, and the greedy approach ensures that we added at least as many edges to the solution as $e(H^{\star}\setminus S,H^{\star}\cap S)$.
@@
2Let $U$ be the set of nodes outside $S$ of size $\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil$ with highest $e(v,S)$ values. Ties broken arbitrarily.
@@
Consider Algorithm 1. We observe that
$|U|=\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil$ $\geq\frac{\epsilon}{1-\epsilon}|S|$
$\geq\frac{\epsilon}{1-\epsilon}(1-\epsilon)|H^{\star}|$
@@
The inequality $|S|\geq(1-\epsilon)|H^{\star}|$ follows from the first property of $S$. Note that we choose $U\subseteq V\setminus S$ in a way that maximizes $e(U,S)$ among all sets whose size is at most $\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil$. We also showed that $|U|\geq|H^{\star}\setminus S|$. Therefore,
@@
$|S\cup U|$ $=|S|+\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil$
@@
It is straightforward to implement this algorithm in $O(m+n)$ time by iterating over each edge $uv$ and, if $u\notin S$ and $v\in S$ (using a hash table to check membership in $S$), incrementing $e(u,S)$. We then use linear-time selection [^2] In practice, sorting or randomized selection is more practical. to find $U$, i.e., the top $\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil$ nodes with the highest $e(v,S)$ values. ∎
@@
$|S\setminus H|$ $\leq\frac{\sum_{v\in S\setminus H}o(v)^{p}}{(1-\delta)^{p}}\penalty 10000\ .$
@@ 2.2 Learning-Augmented Algorithm for Densest At-Most-$k$-Subgraph
We consider the following algorithm, which first adds $\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil$ nodes $U$ that have the highest number of connections to the predicted solution $S$. Then, starting with $T=U\cup S$, it iteratively removes nodes from $T$ that currently have the lowest degree in $T$ so that the final solution has size exactly $k$ (trimming step).
Input: Graph $G=(V,E)$, $S\subseteq V$ that is a $(1-\epsilon)$ partial solution to $H^{\star}$
@@
2Let $r\leftarrow\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil$;
3
4Let $U\subseteq V\setminus S$ be the $r$ nodes with the largest values of $e(v,S)$;
5
@@
$=\textup{d}(T_{0})\frac{n_{0}-2}{n_{0}-1}\cdot\frac{n_{0}-3}{n_{0}-2}\cdot\frac{n_{0}-2}{n_{0}-3}\cdot\ldots\cdot\frac{n_{0}-(r+1)}{n_{0}-r}.$
@@
$|T_{0}|=|S\cup U|=|S|+\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil\leq(1+\epsilon+\epsilon/(1-\epsilon))|H^{\star}|+1.$
@@
In Algorithm 2, the final output $T$ has $k$ nodes. If no node is removed from $T_{0}$, then $T=T_{0}$ and we are done. Otherwise, applying Lemma 4 with $r=|T_{0}|-k$ yields
@@
In each iteration, the algorithm removes a vertex from the nonempty bucket with minimum index, corresponding to a minimum-degree vertex. When a vertex $v$ is removed, we examine all neighbors $u$ of $v$ that remain in the graph. For each such neighbor, the degree of $u$ decreases by one, and $u$ is moved from bucket $B[\deg(u)]$ to bucket $B[\deg(u)-1]$. Because the buckets are doubly linked lists and each vertex stores a pointer to its list node, each such move can be carried out in $O(1)$ time. Each edge is responsible for at most one degree decrement, and therefore the total time spent updating degrees over the entire execution is $O(m)$.
Let $d$ denote the current minimum degree in an iteration. In the next iteration, the minimum degree vertex can be found by scanning $B[d-1],B[d],B[d+1],\dots$ The overall work to scan these lists is $O(n)$.
@@
3 Experiment
We run all experiments on a machine with Ubuntu 24, an AMD Ryzen 5 5600X3D CPU, and 32 GB of RAM. For the densest at-most-$k$-subgraph experiment, we generate labeled data using Julia. The rest of our experiments are done in Python.
3.1 Learning-Augmented Densest Subgraph
We conduct experiments to evaluate the effectiveness of our method. The dataset we use consists of ego-networks of Twitch users who participated in the partnership program in April 2018, as described in [37]. It contains over 127,000 graphs. Each node represents a user, and each edge represents a friendship.
We restrict our experiment to graphs with at least 100 edges. 275 graphs are used to train a random forest predictor. Since the nodes lack attributes, we follow standard practice in graph learning and use local structural features. Specifically, each node is characterized by its degree, the average degree of its neighbors, and the total number of nodes in its graph. These features can be computed for all nodes in $O(m)$ time, which matches the linear time complexity of our algorithm. The label is 1 if a node belongs to the densest subgraph, and 0 otherwise.
To generate ground truth labels for nodes in the training graphs, we compute the exact densest subgraphs using Goldberg’s max-flow-based algorithm, implemented by Tsourakakis [42]. We use a random forest with 10 decision trees as the classifier.
In summary, each data point corresponds to a node, described by its degree, the average degree of its neighbors, and the total number of nodes in its graph. The label indicates whether the node belongs to the densest subgraph. An example densest subgraph is shown in Figure 3.
Figure 3: An example densest subgraph in a Twitch ego-network.
Despite using only degree and average neighbor degree as features, the model performs reasonably well. It achieves a recall 89% of the densest subgraph. See Table 1.
Hereinafter, we refer to the subgraph consisting of nodes predicted by the classifier to be in the densest subgraph as the predictor’s output. Our goal is to compare the output of Algorithm 1 with that of the predictor and Charikar’s greedy algorithm. The primary metric is the output density (i.e., the density of the output subgraph).
Class Precision Recall F1-score Support
0 0.32 0.12 0.17 715
1 0.68 0.89 0.77 1540
Accuracy 0.64 2255
Table 1: Classification report of a random forest to predict whether a node belongs to the densest subgraph using only degree, average neighbor degree, and graph size as features. 0 means not being in the densest subgraph. 1 means being in the densest subgraph.
We evaluate our algorithm on 253 test graphs that are not part of the training set. As shown in Figure 4, where each data point is a graph in the test set, the $y$-axis represents the output density of Algorithm 1, which augments the predictor’s output by adding additional nodes. The $x$-axis corresponds to the density of the predictor’s output. In nearly all cases, our algorithm (with $\epsilon=0.2$) improves the density over the predictor’s (unaugmented) output. On average, Algorithm 1 achieves a 36% higher output density than the predictor’s output. Moreover, it also outperforms the predictor’s output on $93\%$ of the test inputs.
We are also interested in comparing our algorithm with the linear-time greedy algorithm by Charikar that yields a 1/2-approximation. The improvement over Charikar’s greedy algorithm is also evident as shown in Figure 4. On average, Algorithm 1 improves the output density by 31% over Charikar’s greedy algorithm. It also outperforms the Charikar’s greedy algorithm on $94\%$ of the test inputs.
The result is summarized in Table 2 and the improvement distribution is shown in Figure 5. We note that Algorithm 1 and Charikar’s algorithm are both linear-time and can be combined by simply returning the better of the two outputs. Nonetheless, our results show that Algorithm 1 outperforms Charikar’s greedy algorithm in the vast majority of cases.
Metric Improvement
vs. Predictor’s Output
Average Density Improvement 36%
Median Density Improvement 33%
Win Rate (per Graph) 93%
vs. Charikar’s Greedy’s Output
Average Density Improvement 31%
Median Density Improvement 32%
Win Rate (per Graph) 94%
Table 2: Improvement metrics of the augmented algorithm compared to baselines for the densest subgraph problem. Win Rate represents the percentage of test cases where the proposed approach yielded higher density.
Figure 4: Comparison of output densities for the densest subgraph problem. First figure compares Algorithm 1 to the predictor’s raw output, while second figure compares it to Charikar’s greedy algorithm.
Figure 5: Percentage improvement in terms of density for the densest subgraph problem.
3.2 Learning-Augmented Densest At-Most-$k$ Subgraph
For the learning-augmented densest at-most-$k$-subgraph experiment, we use the same dataset and features as in the densest subgraph experiment. We restrict our attention to graphs with at least $100$ edges. In our experiment, we choose $k=15$.
As before, the features consist of the degree, the average neighbor degree, and the number of nodes in the graph. The labels are computed using a brute-force approach in Julia.

### v3

@@ Abstract
We study the densest subgraph problem and its NP-hard densest at-most-$k$ subgraph variant through the lens of learning-augmented algorithms. We show that, given a reasonably accurate predictor that estimates whether a node belongs to the solution (e.g., a machine learning classifier), one can design simple linear-time algorithms that achieve a $(1-\epsilon)$approximation. Finally, we present experimental results demonstrating the effectiveness of our methods for the densest at-most-$k$ subgraph problem on real-world graphs.
@@ 1 Introduction
In this work, we study the densest subgraph problem and its NP-hard variant, the densest at-most-$k$ subgraph, within the framework of learning-augmented algorithms. This approach is motivated by the observation that real-world data often exhibit underlying patterns that can be captured using machine learning. We mainly assume access to a predictor (e.g., a machine-learning model) that estimates whether a node belongs to the densest subgraph. Such a predictor is typically obtained by training a model on a dataset for which ground-truth labels are available.
@@
Figure 1: Suppose that this is the densest subgraph and the predictor identifies the nodes $y_{1},\ldots,y_{5}$. Returning $y_{1},\ldots,y_{5}$ will yield an infinitely bad solution with density 0.
Related work.
Exact algorithms for the densest subgraph problem exist using linear programming or maximum flow [17, 30]. For unweighted graphs, Goldberg’s max-flow based algorithm runs in $O(mn\log n)$ time [33]. One can also approximate the densest subgraph up to a $1-\epsilon$ factor using the multiplicative weights method to solve the dual linear program followed by a clever rounding algorithm [44, 7, 28] which runs in $O(m\cdot 1/\epsilon^{2}\cdot\log^{2}n)$ time. There is, however, a simple and practical greedy $1/2$-approximation that can be implemented in $O(m+n)$ time by Charikar [17]; this algorithm keeps removing the node with the lowest degree from the graph and recording the densest subgraph in the process.
There is also a heuristic proposed by Boob et al. [15] called Greedy++; this heuristic was later shown by Chekuri, Quanrud, and Torres [18] to run in $O\left(\frac{m\Delta\log n}{\epsilon^{2}d(H^{\star})}\right)$ time (which is $O(\epsilon^{-2}mn\log n)$ in the worst case). Chekuri, Quanrud, and Torres [18] also provided a flow-based $\tilde{O}(m/\epsilon)$ time algorithm that obtains a $1-\epsilon$ approximation.
One practical issue with the densest subgraph problem is that the densest subgraph is often too large to be useful in many real-world graphs. Instead, one may consider the problem of finding the densest subgraph with at most $k$ nodes. This variant is NP-hard. Andersen [3] showed that an $\alpha$-approximation for this problem implies an $O(\alpha^{2})$-approximation for the densest $k$ subgraph problem. Furthermore, under the Exponential Time Hypothesis (ETH), no polynomial-time constant-factor approximation exists for the densest $k$ subgraph problem [35]. Hence, the densest at-most-$k$-subgraph problem is also likely difficult to approximate. For the densest $k$ subgraph problem, the current best polynomial time approximation is $O(n^{1/4})$ [11].
@@
We first address the issue of uniqueness. Throughout, we let $H^{\star}$ denote a fixed optimal solution, and the predictor predicts membership for $H^{\star}$. This assumption is often unavoidable in this setting, as is standard in the literature [16, 2].
In real-world instances, the optimal solution is typically unique and, moreover, there are usually no competing optimal solutions. That is, any solution whose objective value is close to optimal must substantially overlap with the optimal one. This type of stability phenomenon has been studied extensively in problems such as clustering, independent set, and Max-Cut (see, e.g., [9, 13, 5, 6]).
However, for the densest subgraph problem, it is possible to enforce uniqueness as follows. If $H_{1}^{\star}$ and $H_{2}^{\star}$ are two solutions to the densest subgraph problem, then $H_{1}^{\star}\cup H_{2}^{\star}$ is also a solution to the densest subgraph problem. To see this, let $\text{OPT}=\textup{d}(H_{1}^{\star})=\textup{d}(H_{2}^{\star})$. Note that $e(H_{1}^{\star})=\text{OPT}|H_{1}^{\star}|$, $e(H_{2}^{\star})=\text{OPT}|H_{2}^{\star}|$, and $e(H_{1}^{\star}\cap H_{2}^{\star})\leq\text{OPT}|H_{1}^{\star}\cap H_{2}^{\star}|$.
@@
Hence, we may enforce uniqueness for the densest subgraph problem by letting $H^{\star}$ be the maximal densest subgraph.
@@
This implies that the symmetric difference between $S$ and $H^{\star}$ is small $|H^{\star}\triangle S|\leq 2\epsilon|H^{\star}|$. Intuitively, the two conditions say that the provided partial solution $S$ has most nodes in $H^{\star}$ and few nodes outside $H^{\star}$. In the above definition, one can think of the predictor as a machine learning classifier that predicts whether $v\in H^{\star}$ based on its attributes.
@@
    Given a $(1-\epsilon)$-partial solution $S$ to $H^{\star}$, there exists an $O(n+m)$-time algorithm that finds a subgraph whose density is at least
@@
    Given an $(\epsilon,\ell_{p})$-probabilistic predictor for $H^{\star}$, for some $p>0$, there exists an $O(n+m)$-time algorithm that finds a subgraph whose density is at least
@@
We first address the robustness of the above result. Since one can run both this algorithm and Charikar’s greedy algorithm in linear $O(m+n)$ time and return the better solution, the worst-case approximation guarantee is therefore $1/2$. In terms of consistency, as the predictor gets better (e.g., $\epsilon$ gets smaller), we get an approximation closer to $1$ without sacrificing the running time.
One may argue that efficient algorithms and heuristics already exist for the densest subgraph problem and that the algorithm in Theorem 1 is only interesting from a theoretical perspective. However, we want to point out that it can be extended to the problem of finding the densest at-most-$k$ subgraph. Even for $\epsilon$ that is not too small (say $\epsilon\approx 0.5$), we obtain a non-trivial constant approximation which is generally not possible assuming ETH.
This leads us to the second main result that addresses the NP-hard variant where we want to find the densest subgraph $H^{\star}$ subject to the constraint $|H^{\star}|\leq k$.
@@
    Given a $(1-\epsilon)$-partial solution $S$ to $H^{\star}$, there exists an $O(n+m)$-time algorithm that finds an at-most-$k$ subgraph whose density is at least
@@
    Given an $(\epsilon,\ell_{p})$-probabilistic predictor for $H^{\star}$, for some $p>0$, there exists an $O(n+m)$-time algorithm that finds an at-most-$k$ subgraph whose density is at least
@@
As this problem is NP-hard and it is likely that no constant-factor approximation exists in polynomial time, robustness is not guaranteed (even in the normal setting). As for consistency, as the prediction error $\epsilon$ decreases, we obtain an approximation ratio closer to $1$.
Application motivation.
In many applications, graphs are generated by the same underlying process, so patterns from past instances can help solve new ones more quickly.
Many real-world graphs also evolve continuously as new data arrives. Social networks gain or lose edges, and biological networks reconfigure over time. Detecting dense components in real time therefore requires repeatedly computing dense subgraphs on updated snapshots, which is important for tasks such as identifying emerging clusters. However, exact algorithms, especially for the densest at-most-$k$ subgraph problem, quickly become expensive.
In other settings, we analyze different parts of a large network that share common structure, such as research areas in citation networks, geographic regions in social networks, or pathways in biological networks. These subproblems are not independent: they often exhibit recurring structural patterns. A predictor trained on one region can therefore guide solutions in related regions. Our learning-augmented approach exploits this shared structure to achieve speedups while preserving near-optimal solutions.
@@
Algorithms and their analysis for the densest subgraph and densest at-most-$k$ subgraph are presented in Section 2. Section 3 presents experiments of our algorithm for the densest at-most-$k$ subgraph problem on the Twitch Ego Nets and High-energy Physics Citation Network (HEP-PH) datasets.
@@ 2.1 Learning-Augmented Algorithms for Densest Subgraph
$H^{\star}\setminus S$$H^{\star}\cap S$
Figure 2: The approximation guarantee comes from the observation that the number of edges induced by $H^{\star}\setminus S$ is small, and the greedy approach ensures that we added at least as many edges to the solution as $|E(H^{\star}\setminus S,H^{\star}\cap S)|$.
@@
2Let $U$ be the set of nodes outside $S$ of size $\min\left\{\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil,|V\setminus S|\right\}$ with highest $e(v,S)$ values. Ties broken arbitrarily.
@@
Consider Algorithm 1. Let
$r:=\min\left\{\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil,\ |V\setminus S|\right\}.$
Then $|U|=r$. We claim that $|U|\geq|H^{\star}\setminus S|$. Indeed,
$\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil\geq\frac{\epsilon}{1-\epsilon}|S|$ $\geq\frac{\epsilon}{1-\epsilon}(1-\epsilon)|H^{\star}|$
@@
The inequality $|S|\geq(1-\epsilon)|H^{\star}|$ follows from the first property of $S$. Moreover, since $H^{\star}\setminus S\subseteq V\setminus S$, we also have $|V\setminus S|\geq|H^{\star}\setminus S|$. Thus, $r\geq|H^{\star}\setminus S|$.
Note that we choose $U\subseteq V\setminus S$ in a way that maximizes $e(U,S)$ among all sets whose size is $r\geq|H^{\star}\setminus S|$. Therefore,
@@
$|S\cup U|$ $\leq|S|+\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil$
@@
It is straightforward to implement this algorithm in $O(m+n)$ time by iterating over each edge $uv$ and, if $u\notin S$ and $v\in S$ (using a hash table to check membership in $S$), incrementing $e(u,S)$. We then use linear-time selection to find $U$. ∎
@@
$|S\setminus H|$ $\leq\frac{\sum_{v\in S\setminus H}o(v)^{p}}{(1-\delta)^{p}}~.$
@@ 2.2 Learning-Augmented Algorithm for Densest At-Most-$k$-Subgraph
We consider the following algorithm, which first adds $\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil$ nodes $U$ that have the highest number of connections to the predicted solution $S$. Then, starting with $T=U\cup S$, it iteratively removes nodes from $T$ that currently have the lowest degree in $T$ so that the final solution has at most $k$ (trimming step).
Input: Graph $G=(V,E)$, $S\subseteq V$ that is a $(1-\epsilon)$ partial solution to $H^{\star}$ and $k\in\mathbb{Z}^{+}$
@@
2Let $r\leftarrow\min\left\{\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil,|V\setminus S|\right\}$;
3Let $U\subseteq V\setminus S$ be the $r$ nodes with the largest values of $e(v,S)$;
4
@@
$=\textup{d}(T_{0})\frac{n_{0}-2}{n_{0}-1}\cdot\frac{n_{0}-3}{n_{0}-2}\cdot\frac{n_{0}-4}{n_{0}-3}\cdot\ldots\cdot\frac{n_{0}-(r+1)}{n_{0}-r}.$
@@
$|T_{0}|=|S\cup U|\leq|S|+\left\lceil\frac{\epsilon}{1-\epsilon}|S|\right\rceil\leq(1+\epsilon+\epsilon/(1-\epsilon))|H^{\star}|+1.$
@@
In Algorithm 2, the final output $T$ has at most $k$ nodes. If no node is removed from $T_{0}$, then $T=T_{0}$ and we are done. Otherwise, applying Lemma  4 with $r=|T_{0}|-k$ yields
@@
In each iteration, the algorithm removes a vertex from the nonempty bucket with minimum index, corresponding to a minimum-degree vertex. When a vertex $v$ is removed, we examine all neighbors $u$ of $v$ that remain in the graph. For each such neighbor, the degree of $u$ decreases by one, and $u$ is moved from bucket $B[\deg(u)]$ to bucket $B[\deg(u)-1]$. Because the buckets are doubly linked lists and each vertex stores a pointer to its list node, each such move can be carried out in $O(1)$ time. Each edge is responsible for at most one degree decrement, and therefore the total time spent updating degrees over the entire execution is $O(n+m)$.
Let $d$ denote the current minimum degree. Since $d$ can be increased at most $n$ times and each vertex is deleted exactly once, the overall running time of the algorithm is $O(m+n)$.
@@
3 Experiments
3.1 Experimental Setup
Setup.
We run all experiments on a machine with Linux Mint 22.3, an AMD Ryzen 9900X CPU with 24 threads, and 64 GB of DDR5 RAM. All experiments are implemented in Julia.
Because existing algorithms for the densest subgraph problem are already quite efficient in practice, we focus on the densest at-most-$k$-subgraph problem. We use two datasets: Twitch Ego Nets and the High-Energy Physics Citation Network from [34].
For this problem, we focus on regimes in which $k$ is not too small relative to $n$. When $k$ is very small compared to $n$, the labels become too imbalanced for a predictor to perform well.
Pruning.
In order to train a predictor for the densest at-most-$k$-subgraph problem, we need to compute the exact solution for some training graphs to obtain ground truth labels. We propose some simple pruning rules that reduce the search space.